OpenStudy (anonymous):

On the first four exams, your grades are 76, 91, 60, and 77. You are hoping to earn a C in the course. This will occur if the average of your five exam grades is greater than or equal to 70 and less than 80. What range of grades on the 5th exam will result in earning a C? Please show work and explain how you got your answer. :)

OpenStudy (valpey):

76+91+60+77+x between 70*5 and 80*5

OpenStudy (anonymous):

So since there are five exams total, and each exam has the same weight we can set up an equation that looks something like this: a+b+c+d+e ---------- = G 5 where a represesnts your first score, b represents your second so on so fourth, and G represents your Grade in the class. you can now seperate it up so you have: a/5 + b/5 + c/5 + d/5 + e/5 = G Now, we can apply YOUR numbers. 76/5 + 91/5 + 60/5 + 77/5 + x/5 = G G= x/5 + (76+91+60+77)/5 G= x/5 + (304)/5 The way this function applies to your question is We are looking for the values of X which gives you G greater than or equal to 70 and less than 80. G=x/5 + 60.8 Now we plug in your minimum desired Grade 70=x/5 + 60.8 9.2=x/5 46=x now your maximum desired grade 80=x/5 + 60.8 19.2=x/5 96 = x so \[46\le x < 96\] does that make sense?

OpenStudy (anonymous):

Wow Yes. Thanks for explaining it to me and making it clear instead of just giving me the answer :)

OpenStudy (anonymous):

Your welcome. I don't like doing people's homework for them but I don't mind helping people understand math :). Let me know if you have another question!

OpenStudy (anonymous):

Okay. I already did most of these problems and got them wrong or didnt understand them and I just need someone to show me where i made my mistakes and how to do the problems. So thanks for the help!

OpenStudy (anonymous):

No problem, any time!